A 5 digit no. formed by usng 0,1,2,3,4,5 without repitition the prob. that it is divisible by 6 ?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

a number is divisibe by 3 and 2 then it is divisible by 6

divisibility by 2

digit at unit place must be even

divisibility by 3

 sum of the digit divisible by 3

  so only two combination is possible 1,2,3,4,5   and 0,1,2,4,5 

    for 1,2,3,4,5

    unit place can be filled by 2 ways

    all other 4 places can be filled by 4*3*2*1 ways

 so total ways  = 2*4*3*2*1 = 48 ways

 for  0,1,2,4,5

   0 is at unit place

  all other 4 places can be filled by 4*3*2*1 ways = 24

  0 is not at the unit place

  unit place can be filled by 2 ways

  all other place can be filled 3*3*2*1

 total ways =2* 3*3*2*1 =36

   so  48+ 24+36 = 108  ways

probability = 108/5*5*4*3*2

                     =.18

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin